# Comparative study of the butterfly velocity in holographic QCD models at finite temperature and chemical potential

### v2

@@ 2.2 Shock Wave and Out-of-time-ordered-correlators (OTOCs)
where $V$ and $W$ are two Hermitian and unitary operators, $\braket{\mathcal{O}}\equiv\text{tr}\left[e^{-\beta H}\mathcal{O}\right]/\text{tr}e^{-\beta H}$ and $\beta=1/T$ is the inverse of the temperature. $C(x,t)$ accounts for the effect that the perturbation $W$ at earlier time $-t$ and position $x=0$ has, on the operator $V$ at time $t=0$ and position $x$. We make use of the negative time because it becomes convenient once we switch to the holographic picture. This commutator of two spatially separated operators generally behaves like [12, 13, 14]
@@
The second term in (2.16) is what is called as the out-of-time-ordered correlator (OTOC), and it entails all the non-trivial information of the commutator. OTOC is used to diagnose chaos in a quantum system, as the vanishing of OTOC implies chaotic time evolution of the system. This can be physically understood by considering the following two states
|$\ket{\psi_{1}}$ $=$ $W(0,-t)V(x,0)\ket{\beta}\,,$
|$\ket{\psi_{2}}$ $=$ $V(x,0)W(0,-t)\ket{\beta}\,,$ (2.17)
where, in the context of shockwave analysis, the state |$\ket{\beta}$ would correspond to a thermofield double state. The physical interpretation of the state |$\ket{\psi_{1}}$ is as follows: At time $t=0$ (and spatial point $x$), state $V(x,0)\ket{\beta}$ is prepared and then evolved backward in time via the unitary time evolution ($e^{iHt}$, where $H$ is the Hamiltonian). A small perturbation $W$ is made and the system is evolved forward in time again to $t=0$. If we act with $W$ at a sufficiently early time in the past and if the system is chaotic, it would interfere with the perturbation due to $V$, thereby preventing the operator $V$ from rematerializing at $t=0$. Similarly, the state |$\ket{\psi_{2}}$ can be understood as follows: a perturbation $W$ is made at time $-t$ (and at $x=0$), then the system evolves to time $t=0$ at which the operator $V$ is inserted at a spatial point $x$. Since the operator $W$ scrambles among the degrees of freedom in its local region in chaotic systems, the overlap between the two states |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$ becomes small. As a result, the OTOC vanishes in chaotic systems, causing the commutator in equation (2.16) to grow.
@@ 2.2.1 Kruskal extension and the shockwave analysis
To understand the bulk picture, we start with the two-sided black hole geometry in the bulk. This corresponds to a thermofield double state |$\ket{TFD}$ at the boundary and is given by [18]
@@
where $Z$ is the thermal partition function, {|$\{\ket{n}\}$ is a complete set of energy eigenstates, and $E_{n}$ are the corresponding eigenvalues. $L$ and $R$ label the two copies of the states. To work with the two-sided geometry, it is convenient to write the planar black hole metric in Kruskal coordinates. We consider the following Kruskal coordinate transformation
@@
Substituting (2.23), (2.12) into (2.24), we get the following.
@@
In the AdS/CFT correspondence, an operator $\hat{Q}$ in the boundary corresponds to a field $\phi(x,t)$ in the bulk. Suppose we act with an operator $W(0,-t)$ on |$\ket{TFD}$ in the right boundary. In the two-sided black hole geometry, this corresponds to a particle coming out from the past interior, reaching the boundary at time $-t$, and then falling towards the future interior. The energy of the particle falling into the black hole increases exponentially, and it depends on the temperature of the black hole
@@
here $E_{o}$ is the energy of the particle when it is near the boundary. Since the energy of the particle gets exponentially blue-shifted, a sufficiently earlier perturbation in the boundary, $W(0,-t)$, leads to a non-trivial modification of the bulk geometry. The large energy of the particle leads to a back reaction in the geometry, which simply corresponds to a shockwave geometry [13]. The energy distribution of this perturbation is compressed along the $u$ direction and stretched along the $v$ direction. This is illustrated in Figure 2). Thus, for sufficiently large time $|t|$, the perturbation gets localized along the horizon $u=0$. The change in the stress-energy tensor sources this localized shockwave:
@@
where, the step function $\Theta(u)$ accounts for the fact that the causal future of the particle is affected after it encounters the horizon $u=0$, and thus, there is no effect on the causal past. $\bar{h}(x)$ is a function that characterizes the amount of shift, and it is to be determined by the Einstein equation of motion. [^2] The shift function $\bar{h}(x)$ should not be confused with function $h(z)$ used in the spacetime metric in Eq. (2.1).
@@
$u\,(v=0)$$v\,(u=0)$$\bar{h}(x)$
Figure 2: The Penrose diagram (corresponding to a two-sided black hole geometry) illustrates the impact of a shockwave generated near the horizon of the black hole. This shockwave, shown as a blue line, originates from the perturbation created by the particle $W$, which is inserted at an early enough time. Without the shockwave, the $V$ particle would have emerged from the past interior region and reached the boundary at $t=0$. However, the presence of the shockwave alters its trajectory, introducing a shift parameterized by $\bar{h}(x)$. This shift results in a delay in the appearance of the operator $V$ at the boundary. When the $W$ perturbation is inserted sufficiently early ( $|t|\gtrsim|t_{*}|$ ) the trajectory of $V$ is displaced so significantly that it falls into the future interior, preventing it from reaching the boundary at all.
With this physical picture in mind, we now proceed to explicitly compute the function $\bar{h}(x)$. The shift in the coordinates (2.28) results in the following modification in the metric
@@
where $T_{o}$ is the initial unperturbed stress tensor and the only non-zero component of $\delta T^{shock}$ is given by (2.27). Substituting (2.27) and (2.29) in the Einstein equation and solving for the $uu$ component, we arrive at the following equation
@@
Here, $d$ is the number of boundary spacetime dimensions. For large $x$ ($|x|>>1$), the function $a(x)$ can be replaced by the Dirac delta function, as the solution depends only on the integral of $a(x)$. Solving for large $x$, one gets the following solution for $\bar{h}(x)$:
@@
Here $\partial_{u}\partial_{v}B(0)$ denotes the double derivative of $B$ evaluated at the horizon and $\{A(0),B(0)\}$ denote the functions $\{A,B\}$ evaluated at the horizon. Comparing (2.32) with (2.15), we can read off the butterfly velocity
@@
2.3 Pole-Skipping
This section discusses the pole-skipping (PS) phenomenon to compute the butterfly velocity. The PS phenomenon occurs when lines of poles and zeros of the retarded Green function intersect, leading to specific values of frequency ($\omega_{*}$) and momentum ($q_{*}$) at which the Green function does not yield a unique value. In [75, 76], it was demonstrated that these special points can be obtained by a hydrodynamical effective field theory description and are given by
$\omega_{*}=i\lambda_{L},\hskip 8.5359ptq_{*}=i\frac{\lambda_{L}}{v_{B}}\,.$ (2.36)
where, as usual, $\lambda_{L}$ and $v_{B}$ are the Lyapunov exponent and the butterfly velocity. The above relations can be realized from the two-point energy density correlation function, which shows a special behavior at points (2.36) in Fourier space. Essentially, the retarded Green function of the energy density takes the form
@@ 2.3 Pole-Skipping
which allows us to define pole-skipping point $(\omega_{*},q_{*})$. Therefore, by computing the special points $\omega_{*}$ and $q_{*}$, one can obtain the chaotic information of the system from Eq. (2.36). Note that the pole and zero lines of the green function can meet at multiple points in the complex momentum space $(\omega,q)$. Our interest here is not in computing all those pole-skipping points. Our objective is only to compute the butterfly velocity using (2.36) from the pole-skipping point, which is related to the prominent chaotic behavior of the system.
Note that $G^{R}_{T^{00}T^{00}}$ has an indeterminate form at $(\omega_{*},q_{*})$. Near the point $(\omega_{*},q_{*})$, the Green function can be evaluated by the following limit
@@
Since the limit depends on the slope $\delta q/\delta\omega$ of the curve at the point $(\omega_{*},q_{*})$, via which the pole-skipping point is approached, $G^{R}_{T^{00}T^{00}}$ is not uniquely defined at the point $(\omega_{*},q_{*})$. Thus the Green function is multi-valued at $(\omega_{*},q_{*})$ [77, 78, 3].
@@
As previously stated, the appropriate Green function is determined by the perturbed equations, which are regular at the black hole horizon and subject to the ingoing boundary condition. For this purpose, we consider the following near-horizon expansion of the metric and the matter perturbations
$X(z)=\sum_{n=0}^{\infty}X^{(n)}(z_{h})\frac{(z-z_{h})^{n}}{n!}.$ (2.42)
where $X$ collectively represents the perturbations of metric and matter fields, and the superscript $n$ denotes the $n^{\text{th}}$ order of derivative with respect to $z$. To proceed, we further need to define weight. Following [79, 80], it is defined for any tensor component as the number of upper $z$-indices minus the number of upper $v$-indices, where lower $v$-indices are considered as upper $z$-indices and vice-versa.
In [79], it was established that for any arbitrary diffeomorphism invariant Einstein gravity with matter fields, the first instance of pole-skipping points happens at frequency $\omega=i(p_{0}-1)2\pi T$, where $p_{0}$ is the highest weight. For metric perturbations like in Eq. (2.41), the highest weight is $p_{0}=2$. Accordingly, from Eq. (2.36), it is easy to see that the first pole-skipping point directly corresponds to maximal Lyapunov exponent and chaos bound $\lambda_{L}=2\pi T$. Note that the butterfly velocity obtained from the first pole skipping point for $p_{0}=2$ matches with the one obtained from OTOC [79]. It suggests that the first pole-skipping point can be obtained from the Einstein field equation of the highest weighted perturbation field.
@@
Substituting the near-horizon expansion (2.42) of the various metric and matter perturbations into $\delta E_{v}^{z}=\delta T_{v}^{z}$ and comparing the coefficients of $(z-z_{h})^{n}$, we obtain a set of equations (say $\boldsymbol{\mathcal{S}}_{n}=0$) that puts a non-trivial constraint on $\delta g_{vv},\delta g_{vx},\delta g_{xx}$, and $\delta\psi_{i}$. The equation $\boldsymbol{\mathcal{S}}_{0}=0$ takes the following form
$\begin{split}\left(2q\delta g_{vx}^{(0)}(z_{h})+\omega\delta g_{xx}^{(0)}(z_{h})\right)&\left(2\omega-ig^{\prime}(z_{h})\sqrt{\frac{h(z_{h})}{f(z_{h})}}\right)+\left(2q^{2}-(d-1)\frac{i\omega k^{\prime}(z_{h})}{\sqrt{h(z_{h})f(z_{h})}}\right)\delta g_{vv}^{(0)}(z_{h})\\ &=4k(z_{h})\left(\delta T_{vv}(z_{h})-\frac{1}{\sqrt{h(z_{h})f(z_{h})}}T_{vz}(z_{h})\delta g_{vv}^{(0)}(z_{h})\right)\,,\end{split}$ (2.43)
@@
$\begin{split}\left(2q\delta g_{vx}^{(0)}(z_{h})+\omega\delta g_{xx}^{(0)}(z_{h})\right)\left(2\omega-4i\pi T\right)+\left(2q^{2}-(d-1)i\omega\frac{k^{\prime}(z_{h})}{\sqrt{h(z_{h})f(z_{h})}}\right)\delta g_{vv}^{(0)}(z_{h})=0\,,\end{split}$ (2.45)
where we have used Eq. (2.12). For any general $\omega$ and $q$, the above equation provides a non-trivial constraint relating $\delta g_{vx}^{(0)}(z_{h})$, $\delta g_{xx}^{(0)}(z_{h})$, and $\delta g_{vv}^{(0)}(z_{h})$. When $\omega=2i\pi T$, the above equation puts no constraint on $\delta g_{vx}^{(0)}(z_{h})$ and $\delta g_{xx}^{(0)}(z_{h})$ and takes the form
@@
If the co-efficient of $\delta g_{vv}^{(0)}(z_{h})$ also vanishes, the equation holds trivially. Thus, the equation (2.45) becomes trivial for the following values of $\omega$ and $q$
@@
For the above values of $\omega$ and $q$, there is one less equation to solve for $\delta g_{vx}^{(0)}(z_{h})$, $\delta g_{xx}^{(0)}(z_{h})$, and $\delta g_{vv}^{(0)}(z_{h})$ than for any other general value of $\omega$ and $q$. As a result, there is an extra linearly independent ingoing solution (extra ingoing mode) to the linearized Einstein equation at the horizon. When we choose the ingoing mode at the above value of $\omega$ and $q$, unlike at generic $\omega$ and $q$, the ratio $B/A$ does not have a definite value, and the energy density retarded Green’s function becomes infinitely multi-valued, as having an extra ingoing mode leads to having no constraints on $A$ and $B$. More details on this can be found in [77]. Therefore, $\omega$ and $q$ given by (2.47) constitute the pole-skipping point $(\omega_{*},q_{*})$ in the Fourier space.
Using the pole-skipping points, we arrive at the following expressions for $\lambda_{L}$ and $v_{B}$.
@@
Notice that this is the same expression that we derived in section 2.1 using the entanglement wedge method. Accordingly, we will get the same thermal and chemical potential dependent spectrum for the butterfly velocity from these two methods. Thus, in the next section, where we explicitly compute $v_{B}$ for various holographic QCD models, and all we have to do is to check whether the matter action of the models satisfy $\delta T_{v}^{z}=0$ to see if the pole-skipping method yields the same result as the other methods.
@@ 3 Butterfly velocity in 1RCBH and 2RCBH holographic models
3.0.1 Background
In this section, we consider two top-down holographic models that describe a strongly coupled quantum field theory at finite density and temperature, namely the 1RCBH and 2RCBH models [50, 51]. The Einstein-Maxwell-dilaton (EMD) action in five dimensions for both of these models is given by:

### v3

@@ 2.2 Shock wave and OTOCs
where $V$ and $W$ are two Hermitian and unitary operators, $\braket{\mathcal{O}}\equiv\text{tr}\left[e^{-\beta H}\mathcal{O}\right]/\text{tr}e^{-\beta H}$, and $\beta=1/T$ is the inverse of the temperature. $C(x,t)$ accounts for the effect that the perturbation $W$ at earlier time t-t and position $x=0$ has, on the operator $V$ at time $t=0$ and position $x$. We make use of the negative time because it becomes convenient once we switch to the holographic picture. This commutator of two spatially separated operators generally behaves like [12, 13, 14]
@@
The second term in (2.16) is what is called the out-of-time-ordered correlator (OTOC), and it entails all the nontrivial information of the commutator. The OTOC is used to diagnose chaos in a quantum system, as the vanishing of the OTOC implies chaotic time evolution of the system. This can be physically understood by considering the following two states:
[…]
@@ 2.2.1 Kruskal extension and the shock wave analysis
[…]
@@
[…]
@@
Substituting (2.23) and (2.12) into (2.24), we get the following:
@@
[…]
@@
where $E_{o}$ is the energy of the particle when it is near the boundary. Since the energy of the particle gets exponentially blueshifted, a sufficiently earlier perturbation in the boundary, W(0,t)W(0,-t), leads to a nontrivial modification of the bulk geometry. The large energy of the particle leads to a backreaction in the geometry, which simply corresponds to a shock wave geometry [13]. The energy distribution of this perturbation is compressed along the $u$ direction and stretched along the $v$ direction. This is illustrated in Figure 2. Thus, for sufficiently large time $|t|$, the perturbation gets localized along the horizon $u=0$. The change in the stress-energy tensor sources this localized shock wave:
@@
where the step function $\Theta(u)$ accounts for the fact that the causal future of the particle is affected after it encounters the horizon $u=0$, and thus, there is no effect on the causal past. $\bar{h}(x)$ is a function that characterizes the amount of shift, and it is to be determined by the Einstein equation of motion. [^2] The shift function $\bar{h}(x)$ should not be confused with function $h(z)$ used in the spacetime metric in Eq. (2.1).
@@
[…]
Figure 2: The Penrose diagram (corresponding to a two-sided black hole geometry) illustrates the impact of a shock wave generated near the horizon of the black hole. This shock wave, shown as a blue line, originates from the perturbation created by the particle $W$, which is inserted at an early enough time. Without the shock wave, the $V$ particle would have emerged from the past interior region and reached the boundary at $t=0$. However, the presence of the shock wave alters its trajectory, introducing a shift parametrized by $\bar{h}(x)$. This shift results in a delay in the appearance of the operator $V$ at the boundary. When the $W$ perturbation is inserted sufficiently early ( $|t|\gtrsim|t_{*}|$ ) the trajectory of $V$ is displaced so significantly that it falls into the future interior, preventing it from reaching the boundary at all.
With this physical picture in mind, we now proceed to explicitly compute the function $\bar{h}(x)$. The shift in the coordinates (2.28) results in the following modification in the metric:
@@
where $T_{o}$ is the initial unperturbed stress tensor and the only nonzero component of $\delta T^{shock}$ is given by (2.27). Substituting (2.27) and (2.29) in the Einstein equation and solving for the $uu$ component, we arrive at the following equation:
@@
Here, $d$ is the number of boundary spacetime dimensions. For large $x$ ($|x|>>1$), the function $a(x)$ can be replaced by the Dirac $\delta$ delta function, as the solution depends only on the integral of $a(x)$. Solving for large $x$, one gets the following solution for $\bar{h}(x)$:
@@
Here, $\partial_{u}\partial_{v}B(0)$ denotes the double derivative of $B$ evaluated at the horizon and $\{A(0),B(0)\}$ denote the functions $\{A,B\}$ evaluated at the horizon. Comparing (2.32) with (2.15), we can read off the butterfly velocity
@@
2.3 Pole skipping
This section discusses the pole-skipping (PS) phenomenon to compute the butterfly velocity. The PS phenomenon occurs when lines of poles and zeros of the retarded Green’s function intersect, leading to specific values of frequency ($\omega_{*}$) and momentum ($q_{*}$) at which the Green’s function does not yield a unique value. In [76, 77], it was demonstrated that these special points can be obtained by a hydrodynamical effective field theory description and are given by
$\omega_{*}=i\lambda_{L},\hskip 8.5359ptq_{*}=i\frac{\lambda_{L}}{v_{B}}\,$, (2.36)
where, as usual, $\lambda_{L}$ and $v_{B}$ are the Lyapunov exponent and the butterfly velocity. The above relations can be realized from the two-point energy density correlation function, which shows a special behavior at points (2.36) in Fourier space. Essentially, the retarded Green’s function of the energy density takes the form
@@ 2.3 Pole skipping
which allows us to define the pole-skipping point $(\omega_{*},q_{*})$. Therefore, by computing the special points $\omega_{*}$ and $q_{*}$, one can obtain the chaotic information of the system from Eq. (2.36). Note that the pole and zero lines of the Green’s function can meet at multiple points in the complex momentum space $(\omega,q)$. Our interest here is not in computing all those pole-skipping points. Our objective is only to compute the butterfly velocity using (2.36) from the pole-skipping point, which is related to the prominent chaotic behavior of the system.
Note that $G^{R}_{T^{00}T^{00}}$ has an indeterminate form at $(\omega_{*},q_{*})$. Near the point $(\omega_{*},q_{*})$, the Green’s function can be evaluated by the following limit:
@@
Since the limit depends on the slope $\delta q/\delta\omega$ of the curve at the point $(\omega_{*},q_{*})$, via which the pole-skipping point is approached, $G^{R}_{T^{00}T^{00}}$ is not uniquely defined at the point $(\omega_{*},q_{*})$. Thus, the Green function’s is multivalued at $(\omega_{*},q_{*})$ [78, 79, 3].
@@
As previously stated, the appropriate Green’s function is determined by the perturbed equations, which are regular at the black hole horizon and subject to the ingoing boundary condition. For this purpose, we consider the following near-horizon expansion of the metric and the matter perturbations:
$X(z)=\sum_{n=0}^{\infty}X^{(n)}(z_{h})\frac{(z-z_{h})^{n}}{n!}\,$, (2.42)
where $X$ collectively represents the perturbations of metric and matter fields, and the superscript $n$ denotes the $n{\text{th}}$ order of derivative with respect to $z$. To proceed, we further need to define weight. Following [80, 81], it is defined for any tensor component as the number of upper $z$ indices minus the number of upper $v$ indices, where lower $v$ indices are considered as upper $z$ indices and vice versa.
In [80], it was established that, for any arbitrary diffeomorphism invariant Einstein gravity with matter fields, the first instance of pole-skipping points happens at frequency $\omega=i(p_{0}-1)2\pi T$, where $p_{0}$ is the highest weight. For metric perturbations like in Eq. (2.41), the highest weight is $p_{0}=2$. Accordingly, from Eq. (2.36), it is easy to see that the first pole-skipping point directly corresponds to maximal Lyapunov exponent and chaos bound $\lambda_{L}=2\pi T$. Note that the butterfly velocity obtained from the first pole-skipping point for $p_{0}=2$ matches the one obtained from the OTOC [80]. It suggests that the first pole-skipping point can be obtained from the Einstein field equation of the highest weighted perturbation field.
@@
Substituting the near-horizon expansion (2.42) of the various metric and matter perturbations into $\delta E_{v}^{z}=\delta T_{v}^{z}$ and comparing the coefficients of $(z-z_{h})^{n}$, we obtain a set of equations (say, $\boldsymbol{\mathcal{S}}_{n}=0$) that puts a nontrivial constraint on $\delta g_{vv},\delta g_{vx},\delta g_{xx}$, and $\delta\psi_{i}$. The equation $\boldsymbol{\mathcal{S}}_{0}=0$ takes the following form:
[…]
@@
[…]
where we have used Eq. (2.12). For any general $\omega$ and $q$, the above equation provides a nontrivial constraint relating $\delta g_{vx}^{(0)}(z_{h})$, $\delta g_{xx}^{(0)}(z_{h})$, and $\delta g_{vv}^{(0)}(z_{h})$. When $\omega=2i\pi T$, the above equation puts no constraint on $\delta g_{vx}^{(0)}(z_{h})$ and $\delta g_{xx}^{(0)}(z_{h})$ and takes the form
@@
If the coefficient of $\delta g_{vv}^{(0)}(z_{h})$ also vanishes, the equation holds trivially. Thus, Eq. (2.45) becomes trivial for the following values of $\omega$ and $q$:
@@
For the above values of $\omega$ and $q$, there is one less equation to solve for $\delta g_{vx}^{(0)}(z_{h})$, $\delta g_{xx}^{(0)}(z_{h})$, and $\delta g_{vv}^{(0)}(z_{h})$ than for any other general value of $\omega$ and $q$. As a result, there is an extra linearly independent ingoing solution (extra ingoing mode) to the linearized Einstein equation at the horizon. When we choose the ingoing mode at the above value of $\omega$ and $q$, unlike at generic $\omega$ and $q$, the ratio $B/A$ does not have a definite value, and the energy density retarded Green’s function becomes infinitely multivalued, as having an extra ingoing mode leads to having no constraints on $A$ and $B$. More details on this can be found in [78]. Therefore, $\omega$ and $q$ given by (2.47) constitute the pole-skipping point $(\omega_{*},q_{*})$ in the Fourier space.
Using the pole-skipping points, we arrive at the following expressions for $\lambda_{L}$ and $v_{B}$:
@@
Notice that this is the same expression that we derived in Section 2.1 using the entanglement wedge method. Accordingly, we will get the same thermal- and chemical-potential-dependent spectrum for the butterfly velocity from these two methods. Thus, in the next section, where we explicitly compute $v_{B}$ for various holographic QCD models, and all we have to do is to check whether the matter action of the models satisfies $\delta T_{v}^{z}=0$ to see if the pole-skipping method yields the same result as the other methods.
@@ 3 Butterfly velocity in 1RCBH and 2RCBH holographic models
3.1 Background
In this section, we consider two top-down holographic models that describe a strongly coupled quantum field theory at finite density and temperature, namely, the 1RCBH and 2RCBH models [51, 52]. The Einstein-Maxwell-dilaton (EMD) action in five dimensions for both of these models is given by